# Lagrangian Skeleta and Koszul Duality on Bionic Symplectic Varieties

### v1

@@ 2.6. Hypertoric varieties and Nakajima quiver varieties
Similarly, Nakajima quiver varieties often give rise to bionic symplectic varieties. Indeed, the example of section 2.5 can be realised as a Nakajima quiver varieties associated to the cyclic quiver. Another large class of examples come from framed finite ADE type quivers. Using standard notation, if $\mathbf{v}$ is the dimension vector and $\mathbf{w}$ the framing vector then we require that each weight $\mathbf{w}_{i}\varpi_{i}$ is a minisucle weight for the corresponding simple Lie algebra. In this case, the smooth quiver variety $M(\mathbf{v},\mathbf{w})$ is bionic; see [Kam, Theorem 6.2].
@@ 2.7. Slodowy Slices
If $\mathfrak{g}$ is a reductive Lie algebra and $e\in\mathfrak{g}$ is a nilpotent element, Slodowy [Sl] has defined a natural transverse slice $S_{e}$ at $e$ through its orbit $\mathcal{O}_{e}$. Let $\mathcal{S}_{e}=S_{e}\cap\mathcal{N}$ be the intersection of the slice with the nilpotent cone of $\mathfrak{g}$. Pulling back the Springer resolution to this slice one obtains a resolution of singularities of $\pi_{e}\colon\Lambda_{e}\to\mathcal{S}_{e}$, which is compatible with the natural symplectic forms on the smooth locus of $\mathcal{S}_{e}$ and $\Lambda_{e}$ so that $\pi_{e}$ is a symplectic resolution.
There is a natural $\mathbb{G}_{m}$-action on $\mathcal{S}_{e}$ and $\Lambda_{e}$ (for which $\pi_{e}$ is equivariant) which contracts $\mathcal{S}_{e}$ to the element $e$: Indeed if $\mathfrak{s}$ is an $\mathfrak{sl}_{2}$-triple containing $e$, then there is a one-parameter subgroup $\lambda$ of $G$ parametrizing the maximal torus of the associated $\text{SL}_{2}$ in $G$. This acts with weight $2$ on $e$, hence if we define $\rho(t)(x)=t^{-2}\lambda(t)(x)$ ($x\in\mathfrak{g})$, then $\rho$ fixes $e$. A tangent space computation shows that it contracts the slice $S_{e}$ to $e$, and clearly it lifts to an action on $\Lambda_{e}$. Since the action is compatible with the map $\pi_{e}$ and $\pi^{-1}(e)$ is proper, every $x\in\Lambda_{e}$ has a limit under the action of $\rho(\mathbb{G}_{m})$. Moreover, the action of $\lambda$ on $T^{*}(\mathcal{B})$ is Hamiltonian, and so the symplectic form $\omega$ on $\Lambda_{e}$ is only altered by scaling of the fibre direction. It follow $\rho(\mathbb{G}_{m})$ has weight $2$ on $\omega$, so that $\rho$ is an elliptic action.
Not all $\Lambda_{e}$ appear to be bionic symplectic varieties however, but a large class of them are, as the following shows: Recall that we say a nilpotent $e\in\mathfrak{g}$ is of standard Levi type if there is a Levi subalgebra $\mathfrak{g}_{0}$ containing $e$ in which $e$ is regular.
@@
Suppose that $e$ is of standard Levi type. Then $\Lambda_{e}$ is a bionic symplectic variety.
@@
By the previous paragraph, it suffices to construct a Hamiltonian $\mathbb{G}_{m}$ action $\rho\colon\mathbb{G}_{m}\to\text{Aut}(\Lambda_{e})$ which commutes with the Slodowy $\mathbb{G}_{m}$-action and has finitely many fixed points.
@@
Now since $e\in\mathfrak{g}_{0}$ is regular, there is a unique Borel subalgebra $\mathfrak{b}_{0}\subseteq\mathfrak{g}_{0}$ containing $e$. But if $\mathcal{B}$ and $\mathcal{B}_{0}$ denote the flag varieties of $\mathfrak{g}$ and $\mathfrak{g}_{0}$ respectively, then $\mathcal{B}^{\lambda(\mathbb{G}_{m})}$ is isomorphic to $|W/W_{0}|$ copies of $\mathfrak{B}_{0}$, where on each component the map $\mathfrak{b}\mapsto\mathfrak{b}\cap\mathfrak{g}_{0}$ yields the isomorphism. Moreover, if $x\in\mathfrak{g}_{0}\cap\mathcal{S}_{e}$, then since $e$ is regular in $G_{0}$, the orbit of $x$ under $G_{0}$ (the Levi subgroup associated to $\mathfrak{g}_{0}$) lies in the closure of the orbit of $e$, and hence the closure of the $G$-orbit of $e$ intersects (and hence contains) the orbit of $x$. Since $S_{e}$ is a transverse slice through the $G$-orbit of $e$, this is possible if and only if $x=e$.
Now suppose that $(x,\mathfrak{b})\in\Lambda_{e}$ is fixed by $\lambda(\mathbb{G}_{m})$, then $x\in\mathfrak{g}_{0}\cap\mathcal{S}_{e}=\{e\}$ and $\mathfrak{b}\cap\mathfrak{g}_{0}=\mathfrak{b}_{0}$, hence there are exactly $|W/W_{0}|$ fixed points for the action of $\lambda(\mathbb{G}_{m})$. It follows $\Lambda_{e}$ is bionic as required. ∎
@@
The condition that $e$ be a nilpotent element of regular Levi type may also be necessary: if the one-parameter subgroup $\lambda$ arises from a one-parameter subgroup of $T\times\mathbb{G}_{m}$ where $T$ is a maximal torus of $G$ and $\mathbb{G}_{m}$ acts by homotheties on $\mathfrak{g}$ then the action will be Hamiltonian if and only if the one-parameter subgroup has image in $T$. But then the condition that $\lambda(\mathbb{G}_{m})$ has finitely many fixed points ensures that $\pi_{e}^{-1}(e)$ is a finite set, which implies that $e$ is regular in the Levi subalgebra given by $\lambda$. Thus if $\text{Aut}(\Lambda_{e})$ has maximal torus $T\times\mathbb{G}_{m}$ (that is, if there are no other one-parameter subgroups up to conjugacy) we obtain the converse to the lemma.
@@
3. ${\mathcal{W}}$-Modules Supported on $\mathsf{\Lambda}$
@@ 3.1. Background on ${\mathcal{W}}$-Algebras and Quasicoherent ${\mathcal{W}}$-Modules
Given an open immersion $j:U\rightarrow\mathfrak{X}$, there is an exact, full, continuous, and essentially surjective restriction functor
@@
There exist adjoint pairs $(j^{*},j_{*})$ and $(i_{*},\mathbb{R}i^{!})$
${D({\mathcal{W}}_{S}\text{-}{\mathsf{qcoh}})}$${D({\mathcal{W}}_{\mathfrak{X}}\text{-}{\mathsf{qcoh}})}$${D({\mathcal{W}}_{U}\text{-}{\mathsf{qcoh}})}$$\scriptstyle{i_{*}}$$\scriptstyle{\mathbb{R}i^{!}}$$\scriptstyle{j^{*}}$$\scriptstyle{j_{*}}$
@@
(3.2) $i_{*}\circ\mathbb{R}i^{!}\to\operatorname{Id}\to j_{*}j^{*}\stackrel{{\scriptstyle[1]}}{{\longrightarrow}}.$
@@
The fact that the pair $(j^{*},j_{*})$ exists is [BDMN, Corollary 5.9], and that the pair $(i_{*},\mathbb{R}i^{!})$ exists is Proposition 5.1 and Lemma 5.3 of [BDMN]. Finally, it is shown in (5.1) of [BDMN] that $j^{*}$ identifies $D({\mathcal{W}}_{U}\text{-}{\mathsf{qcoh}})$ with $D({\mathcal{W}}\text{-}{\mathsf{qcoh}})/D_{C}({\mathcal{W}}\text{-}{\mathsf{qcoh}})$. Therefore, the other claim is standard, see for instance [KS1, Ex. 10.15]. ∎
3.2. Holonomic Objects and Objects Supported on $\mathsf{\Lambda}$
@@
This paper is concerned mainly with those $\mathsf{T}$-equivariant holonomic objects supported on $\mathsf{\Lambda}$. Consider an open subset $U=\mathfrak{X}\smallsetminus K_{>i}$ for a closed union $K_{>i}$ of coisotropic cells. Then $C_{i}$ is closed in $U$ and $\mathsf{\Lambda}_{i}$ is closed in $C_{i}$. Recall that $\overline{\mathsf{\Lambda}}_{i}=\pi_{i}(\mathsf{\Lambda}_{i})$ where $\pi_{i}:C_{i}\rightarrow S_{i}$ is the coisotropic reduction. In [BDMN] it is shown that if $\mathcal{W}_{\mathfrak{X}}$ is a $\mathbb{G}_{m}$-equivariant quantization of $\mathfrak{X}$ then each $S_{i}$ has a natural $\mathbb{G}_{m}$-equivariant quantization $\mathcal{W}_{S_{i}}$, and there is a reduction functor $\mathbb{H}_{i}$ giving an equivalence between the equivariant modules for $\mathcal{W}_{S_{i}}$ and the subquotient category of modules supported on $C_{i}$ (in both the abelian and dg contexts).
@@
    There is a simple $\mathsf{T}$-equivariant holonomic object $\delta_{\mathsf{\Lambda}_{i}}$ of ${\mathcal{W}}_{U}\text{-}{\mathsf{good}}$ supported on $\mathsf{\Lambda}_{i}$.
@@
    Such an object $\delta_{\mathsf{\Lambda}_{i}}$ is unique up to twisting by a character of the Hamiltonian ${\mathbb{G}}_{m}$-action.
@@
    every $\mathsf{T}$-equivariant holonomic object of ${\mathcal{W}}_{U}\text{-}{\mathsf{good}}$ supported on $\mathsf{\Lambda}_{i}$ is isomorphic to a finite direct sum of twists (for the Hamiltonian ${\mathbb{G}}_{m}$-action) of copies of $\delta_{\mathsf{\Lambda}_{i}}$.
@@
Let ${\mathcal{W}}_{S_{i}}$ denote the ${\mathbb{G}}_{m}$-equivariant ${\mathcal{W}}$-algebra on $S_{i}$. By Theorem 1.3 of [BDMN], checking that each statement of the proposition holds on $\mathsf{\Lambda}_{i}\subset U$ is equivalent to checking the corresponding statement on $\overline{\mathsf{\Lambda}}_{i}\subset S_{i}$. By [BDMN], the category of good ${\mathcal{W}}_{S_{i}}$-modules is equivalent to the category of coherent ${\mathcal{D}}({\mathbb{A}}^{d_{i}})$-modules where $2d_{i}=\dim(S_{i})$. The additional equivariance with respect to the Hamiltonian action on $\mathfrak{X}$ translates to monodromicity of ${\mathcal{D}}({\mathbb{A}}^{d_{i}})$-modules. Hence the claim is that there is a unique simple monodromic ${\mathcal{D}}({\mathbb{A}}^{d_{i}})$-module supported (microlocally) on the “attracting locus” in $T^{*}V$ where $V$ is a ${\mathbb{G}}_{m}$-representation with positive weights. The attracting locus is the zero section $V\subset T^{*}V$ and we need to show that there is a unique simple ${\mathbb{G}}_{m}$-monodromic local system on a ${\mathbb{G}}_{m}$-representation with positive weights.
Let $\mathcal{D}=\mathcal{D}(\mathbb{A}^{d_{i}})$ denote the Weyl algebra equipped with the filtration $F_{\bullet}$ induced by the elliptic action on $\mathcal{W}_{S_{i}}$. The Hamiltonian action $\mathsf{H}$ induces a $\mathbb{Z}$-grading $\mathcal{D}=\bigoplus_{n\in\mathbb{Z}}\mathcal{D}_{n}$ where $\deg x_{i}>0$ and $\deg\partial_{i}=-\deg x_{i}$. Note that $\mathcal{D}_{0}={\mathbb{C}}\langle x_{i}\partial_{i}\rangle$ is a commutative subalgebra. Since the elliptic and Hamiltonian actions commute, the filtration and grading are compatible. Moreover, since $F_{k}\mathcal{D}$ is finite dimensional, so is $F_{k}\mathcal{D}\cap\mathcal{D}_{n}$ for every $n$. Suppose that $M$ is a simple monodromic $\mathcal{D}$-module supported on $V$. Let $M=\bigoplus_{n\in\mathbb{Z}}M_{n}$ be the grading on $M$ given by the Hamiltonian $\mathbb{G}_{m}$-action. Since $M$ is supported on $V$, we claim that $M_{n}=0$ for all $n$ sufficiently large and negative. To see this, note that since the associated graded of $M$ (with respect to the filtration $F_{\bullet}$) is supported on $V$, the generators of $\text{gr}_{F}\mathcal{D}$ with negative grading must act nilpotently. Since $\mathrm{gr}_{F}M$ is finitely generated, the result follows.

### v2

@@ 2.6. Hypertoric varieties and Nakajima quiver varieties
Similarly, Nakajima quiver varieties often give rise to bionic symplectic varieties. Indeed, the example of Section 2.5 can be realised as a Nakajima quiver variety associated to the cyclic quiver. Another large class of examples come from framed finite ADE type quivers. Using standard notation, if $\mathbf{v}$ is the dimension vector and $\mathbf{w}$ the framing vector then we require each weight $\mathbf{w}_{i}\varpi_{i}$ to be a minisucle weight for the corresponding simple Lie algebra. In this case, the smooth quiver variety $M(\mathbf{v},\mathbf{w})$ is bionic; see [Kam, Theorem 6.2].
@@ 2.7. Slodowy Slices
If $\mathfrak{g}$ is a reductive Lie algebra and $e\in\mathfrak{g}$ is a nilpotent element, Slodowy [Sl] has defined a natural transverse slice $S_{e}$ at $e$ through its orbit $\mathcal{O}_{e}$. Let $\mathcal{S}_{e}=S_{e}\cap\mathcal{N}$ be the intersection of the slice with the nilpotent cone of $\mathfrak{g}$. Pulling back the Springer resolution to this slice one obtains a resolution of singularities of $\pi_{e}\colon\widetilde{\mathcal{S}}_{e}\to\mathcal{S}_{e}$, which is compatible with the natural symplectic forms on the smooth locus of $\mathcal{S}_{e}$ and $\widetilde{\mathcal{S}}_{e}$ so that $\pi_{e}$ is a symplectic resolution.
There is a natural $\mathbb{G}_{m}$-action on $\mathcal{S}_{e}$ and $\widetilde{\mathcal{S}}_{e}$ (for which $\pi_{e}$ is equivariant) which contracts $\mathcal{S}_{e}$ to the element $e$: Indeed if $\mathfrak{s}$ is an $\mathfrak{sl}_{2}$-triple containing $e$, then there is a one-parameter subgroup $\lambda$ of $G$ parametrizing the maximal torus of the associated $\text{SL}_{2}$ in $G$. This acts with weight $2$ on $e$, hence if we define $\rho(t)(x)=t^{-2}\lambda(t)(x)$ ($x\in\mathfrak{g})$, then $\rho$ fixes $e$. A tangent space computation shows that it contracts the slice $S_{e}$ to $e$, and clearly it lifts to an action on $\widetilde{\mathcal{S}}_{e}$. Since the action is compatible with the map $\pi_{e}$ and $\pi^{-1}(e)$ is proper, every $x\in\widetilde{\mathcal{S}}_{e}$ has a limit under the action of $\rho(\mathbb{G}_{m})$. Moreover, the action of $\lambda$ on $T^{*}(\mathcal{B})$ is Hamiltonian, and so the symplectic form $\omega$ on $\widetilde{\mathcal{S}}_{e}$ is only altered by scaling of the fibre direction. It follows that $\rho(\mathbb{G}_{m})$ has weight $2$ on $\omega$, so that $\rho$ is an elliptic action. There is a natural $\mathbb{G}_{m}$-action on $\mathcal{S}_{e}$ and $\widetilde{\mathcal{S}}_{e}$ (for which $\pi_{e}$ is equivariant) which contracts $\mathcal{S}_{e}$ to the element $e$: Indeed if $\mathfrak{s}$ is an $\mathfrak{sl}_{2}$-triple containing $e$, then there is a one-parameter subgroup $\lambda$ of $G$ parametrizing the maximal torus of the associated $\text{SL}_{2}$ in $G$. This acts with weight $2$ on $e$, hence if we define $\rho(t)(x)=t^{-2}\lambda(t)(x)$ ($x\in\mathfrak{g})$, then $\rho$ fixes $e$. A tangent space computation shows that it contracts the slice $S_{e}$ to $e$, and clearly it lifts to an action on $\widetilde{\mathcal{S}}_{e}$. Since the action is compatible with the map $\pi_{e}$ and $\pi^{-1}(e)$ is proper, every $x\in\widetilde{\mathcal{S}}_{e}$ has a limit under the action of $\rho(\mathbb{G}_{m})$. Moreover, the action of $\lambda$ on $T^{*}(\mathcal{B})$ is Hamiltonian, and so the symplectic form $\omega$ on $\widetilde{\mathcal{S}}_{e}$ is only altered by scaling of the fibre direction. It follows that $\rho(\mathbb{G}_{m})$ has weight $2$ on $\omega$, so that $\rho$ is an elliptic action.
Not all $\widetilde{\mathcal{S}}_{e}$ appear to be bionic symplectic varieties however, but a large class of them are, as the following shows: Recall that we say a nilpotent $e\in\mathfrak{g}$ is of standard Levi type if there is a Levi subalgebra $\mathfrak{g}_{0}$ containing $e$ in which $e$ is regular.
@@
Suppose that $e$ is of standard Levi type. Then $\widetilde{\mathcal{S}}_{e}$ is a bionic symplectic variety.
@@
By the previous paragraph, it suffices to construct a Hamiltonian $\mathbb{G}_{m}$ action $\rho\colon\mathbb{G}_{m}\to\text{Aut}(\widetilde{\mathcal{S}}_{e})$ which commutes with the Slodowy $\mathbb{G}_{m}$-action and has finitely many fixed points.
@@
Now since $e\in\mathfrak{g}_{0}$ is regular, there is a unique Borel subalgebra $\mathfrak{b}_{0}\subseteq\mathfrak{g}_{0}$ containing $e$. But if $\mathcal{B}$ and $\mathcal{B}_{0}$ denote the flag varieties of $\mathfrak{g}$ and $\mathfrak{g}_{0}$ respectively, then $\mathcal{B}^{\lambda(\mathbb{G}_{m})}$ is isomorphic to $|W/W_{0}|$ copies of $\mathcal{B}_{0}$, where on each component the map $\mathfrak{b}\mapsto\mathfrak{b}\cap\mathfrak{g}_{0}$ yields the isomorphism. Moreover, if $x\in\mathfrak{g}_{0}\cap\mathcal{S}_{e}$, then since $e$ is regular in $\mathfrak{g}_{0}$, the orbit of $x$ under $G_{0}$ (the Levi subgroup associated to $\mathfrak{g}_{0}$) lies in the closure of the orbit of $e$, and hence the closure of the $G$-orbit of $e$ intersects (and hence contains) the orbit of $x$. Since $S_{e}$ is a transverse slice through the $G$-orbit of $e$, this is possible if and only if $x=e$.
Now suppose that $(x,\mathfrak{b})\in\widetilde{\mathcal{S}}_{e}$ is fixed by $\lambda(\mathbb{G}_{m})$, then $x\in\mathfrak{g}_{0}\cap\mathcal{S}_{e}=\{e\}$ and $\mathfrak{b}\cap\mathfrak{g}_{0}=\mathfrak{b}_{0}$, hence there are exactly $|W/W_{0}|$ fixed points for the action of $\lambda(\mathbb{G}_{m})$. It follows that $\widetilde{\mathcal{S}}_{e}$ is bionic as required. ∎
@@
The condition that $e$ be a nilpotent element of regular Levi type may also be necessary: if the one-parameter subgroup $\lambda$ arises from a one-parameter subgroup of $T\times\mathbb{G}_{m}$, where $T$ is a maximal torus of $G$ and $\mathbb{G}_{m}$ acts by homotheties on $\mathfrak{g}$, then the action will be Hamiltonian if and only if the one-parameter subgroup has image in $T$. But then the condition that $\lambda(\mathbb{G}_{m})$ has finitely many fixed points ensures that $\pi_{e}^{-1}(e)$ is a finite set, which implies that $e$ is regular in the Levi subalgebra given by $\lambda$. Thus if $\text{Aut}(\widetilde{\mathcal{S}}_{e})$ has maximal torus $T\times\mathbb{G}_{m}$ (that is, if there are no other one-parameter subgroups up to conjugacy) we obtain the converse to the lemma.
@@
3. ${\mathcal{W}}$-Modules Supported on $\Lambda^{c}$
@@ 3.1. Background on ${\mathcal{W}}$-Algebras and Quasicoherent ${\mathcal{W}}$-Modules
Given an open immersion $j:U\rightarrow\mathfrak{X}$ from a $\mathbb{G}_{m}$-invariant open set $U$, there is an exact, full, continuous, and essentially surjective restriction functor
@@
There exist adjoint pairs $(j^{*},\mathbb{R}j_{*})$ and $(i_{*},\mathbb{R}i^{!})$
${D({\mathcal{W}}_{S}\text{-}{\mathsf{qcoh}})}$${D({\mathcal{W}}_{\mathfrak{X}}\text{-}{\mathsf{qcoh}})}$${D({\mathcal{W}}_{U}\text{-}{\mathsf{qcoh}})}$$\scriptstyle{i_{*}}$$\scriptstyle{\mathbb{R}i^{!}}$$\scriptstyle{j^{*}}$$\scriptstyle{\mathbb{R}j_{*}}$
@@
(3.2) $i_{*}\circ\mathbb{R}i^{!}\to\operatorname{Id}\to\mathbb{R}j_{*}j^{*}\stackrel{{\scriptstyle[1]}}{{\longrightarrow}}.$
@@
The fact that the pair $(j^{*},\mathbb{R}j_{*})$ exists is [BDMN, Corollary 5.9], and that the pair $(i_{*},\mathbb{R}i^{!})$ exists is Proposition 5.1 and Lemma 5.3 of [BDMN]. Finally, it is shown in (5.1) of [BDMN] that $j^{*}$ identifies $D({\mathcal{W}}_{U}\text{-}{\mathsf{qcoh}})$ with $D({\mathcal{W}}\text{-}{\mathsf{qcoh}})/D_{C}({\mathcal{W}}\text{-}{\mathsf{qcoh}})$. Therefore, the other claim is standard, see for instance [KS1, Ex. 10.15]. ∎
3.2. Holonomic Objects and Objects Supported on $\Lambda^{c}$
@@
This paper is concerned mainly with those $\mathsf{T}$-equivariant holonomic objects which are supported on $\Lambda^{c}$. Consider an open subset $U=\mathfrak{X}\smallsetminus K_{>i}$ for a closed union $K_{>i}$ of coisotropic cells. Then $C_{i}$ is closed in $U$ and $\Lambda^{c}_{i}$ is closed in $C_{i}$. Recall that $\overline{\Lambda^{c}}_{i}=\pi_{i}(\Lambda^{c}_{i})$ where $\pi_{i}:C_{i}\rightarrow S_{i}$ is the coisotropic reduction. In [BDMN] it is shown that if $\mathcal{W}_{\mathfrak{X}}$ is a $\mathbb{G}_{m}$-equivariant quantization of $\mathfrak{X}$ then each $S_{i}$ has a natural $\mathbb{G}_{m}$-equivariant quantization $\mathcal{W}_{S_{i}}$ and there is a reduction functor $\mathbb{H}_{i}$ giving an equivalence
$\mathbb{H}_{i}\colon{\mathcal{W}}_{U}\text{-}{\mathsf{good}}_{C_{i}}{\;\stackrel{{{}_{\sim}}}{{\to}}\;}{\mathcal{W}}_{S_{i}}\text{-}{\mathsf{good}}$
between the category of good modules supported on $C_{i}$ and the category of good modules for $\mathcal{W}_{S_{i}}$. The analogous statement holds in the dg context as well.
@@
    There is a simple $\mathsf{T}$-equivariant holonomic object $\delta_{\Lambda^{c}_{i}}$ of ${\mathcal{W}}_{U}\text{-}{\mathsf{good}}$ supported on $\Lambda^{c}_{i}$.
@@
    Such an object $\delta_{\Lambda^{c}_{i}}$ is unique up to twisting by a character of the Hamiltonian ${\mathbb{G}}_{m}$-action.
@@
    every $\mathsf{T}$-equivariant holonomic object of ${\mathcal{W}}_{U}\text{-}{\mathsf{good}}$ supported on $\Lambda^{c}_{i}$ is isomorphic to a finite direct sum of twists (for the Hamiltonian ${\mathbb{G}}_{m}$-action) of copies of $\delta_{\Lambda^{c}_{i}}$.
@@
Let ${\mathcal{W}}_{S_{i}}$ denote the ${\mathbb{G}}_{m}$-equivariant ${\mathcal{W}}$-algebra on $S_{i}$. By Theorem 1.5 of [BDMN], checking that each statement of the proposition holds on $\Lambda^{c}_{i}\subset U$ is equivalent to checking the corresponding statement on $\overline{\Lambda^{c}}_{i}\subset S_{i}$. By [BDMN, Corollary 1.6], the category of good ${\mathcal{W}}_{S_{i}}$-modules is equivalent to the category of coherent ${\mathcal{D}}({\mathbb{A}}^{d_{i}})$-modules where $2d_{i}=\dim(S_{i})$. The additional equivariance with respect to the Hamiltonian action on $\mathfrak{X}$ translates to monodromicity of ${\mathcal{D}}({\mathbb{A}}^{d_{i}})$-modules. Hence the claim is that there is a unique simple monodromic ${\mathcal{D}}({\mathbb{A}}^{d_{i}})$-module supported (microlocally) on the “attracting locus” in $T^{*}V$ where $V$ is a ${\mathbb{G}}_{m}$-representation with positive weights. The attracting locus is the zero section $V\subset T^{*}V$ and we need to show that there is a unique simple ${\mathbb{G}}_{m}$-monodromic local system on a ${\mathbb{G}}_{m}$-representation with positive weights.
Let $\mathcal{D}=\mathcal{D}(\mathbb{A}^{d_{i}})$ denote the Weyl algebra equipped with the filtration $F_{\bullet}$ induced by the elliptic action on $\mathcal{W}_{S_{i}}$. The Hamiltonian action $\mathsf{H}$ induces a $\mathbb{Z}$-grading $\mathcal{D}=\bigoplus_{n\in\mathbb{Z}}\mathcal{D}_{n}$ where $\deg x_{i}>0$ and $\deg\partial_{i}=-\deg x_{i}$. Note that $\mathcal{D}_{0}={\mathbb{C}}\langle x_{i}\partial_{i}\rangle$ is a commutative subalgebra. Since the elliptic and Hamiltonian actions commute, the filtration and grading are compatible. Moreover, since $F_{k}\mathcal{D}$ is finite dimensional, so is $F_{k}\mathcal{D}\cap\mathcal{D}_{n}$ for every $n$. Suppose that $M$ is a simple monodromic $\mathcal{D}$-module supported on $V$. Let $M=\bigoplus_{n\in\mathbb{Z}}M_{n}$ be the grading on $M$ given by the Hamiltonian $\mathbb{G}_{m}$-action. Since $M$ is supported on $V$, we claim that $M_{n}=0$ for all $n$ sufficiently negative. To see this, note that since the associated graded of $M$ (with respect to the filtration $F_{\bullet}$) is supported on $V$, the generators of $\text{gr}_{F}\mathcal{D}$ with negative grading must act nilpotently. Since $\mathrm{gr}_{F}M$ is finitely generated, the result follows.